# Gamma-Ray Burst Light Curve Reconstruction: A Comparative Machine and Deep Learning Analysis

### v4

@@ Abstract
Gamma-Ray Bursts (GRBs), observed at large redshifts, are probes of the evolution of the Universe and can be used as cosmological tools. To this end, we need tight (with small dispersion) correlations among key parameters. To reduce such a dispersion, we will mitigate gaps in light curves (LCs), including the plateau region, key to building the two-dimensional Dainotti relation between the end time of plateau emission ($T_{a}$) and its luminosity ($L_{a}$). We reconstruct LCs using nine models: Multi-Layer Perceptron (MLP), Bi-Mamba, Fourier Transform, Gaussian Process–Random Forest Hybrid (GP-RF), Bidirectional Long Short-Term Memory (Bi-LSTM), Conditional GAN (CGAN), SARIMAX-based Kalman filter, Kolmogorov–Arnold Networks (KANs), and Attention U-Net. These methods are compared to the Willingale model (W07) over a sample of 545 GRBs. MLP and Bi-Mamba outperform other methods, with MLP reducing the plateau parameter uncertainties by 25.9% for $\log T_{a}$, 28.6% for $\log F_{a}$, and 37.7% for $\alpha$ (the post-plateau slope in the W07 model), achieving the lowest 5-fold cross validation (CV) mean squared error (MSE) of 0.0275. Bi-Mamba achieved the lowest uncertainty of parameters, a 33.3% reduction in $\log T_{a}$, a 33.6% reduction in $\log F_{a}$ and a 41.9% reduction in $\alpha$, but with a higher MSE of 0.130. Bi-Mamba brings the lowest outlier percentage for $\log T_{a}$ and $\log F_{a}$ (2.70%), while MLP carry $\alpha$ outliers to 0.900%. The other methods yield MSE values ranging from 0.0339 to 0.174. These improvements in parameter precision are needed to use GRBs as standard candles, investigate theoretical models, and predict GRB redshifts through machine learning.
@@ 1 Introduction
GRBs are brief and highly luminous astrophysical phenomena detectable at remarkable distances (for review, see Kumar & Zhang, 2015), with observations extending up to redshift $z=9.4$ (Cucchiara et al., 2011). This property makes GRBs exceptional cosmological tools for probing the universe’s early evolution. A thorough analysis of GRBs also yields essential information about Population III stars, the first generation formed during the epoch of re-ionization. GRB emission is usually observed in two episodes: the prompt and the afterglow. The prompt phase is interpreted by internal shell collisions or magnetic reconnection and is characterized by the short duration and high energy. Hence, this is detected predominantly in $\gamma$ rays and X-rays and occasionally in optical wavelengths (Vestrand et al., 2005; Blake et al., 2005; Beskin et al., 2010; Gorbovskoy et al., 2012; Vestrand et al., 2014). During the afterglow phase, which follows the primary episode, the relativistic jet impacts the circumstellar environment, transferring part of its energy. The afterglow is characterized by a long duration and energies that span a broad spectrum, including X-ray, optical, and sometimes radio bands (Costa et al., 1997; van Paradijs et al., 1997; Piro et al., 1998; Gehrels et al., 2009).
@@
The Neil Gehrels Swift Observatory (Swift, Gehrels et al. (2004)) is vital for detecting the temporal properties of GRBs. The Swift Burst Alert Telescope (BAT), which operates in the range of 15-150 keV (Barthelmy et al., 2005), plays a crucial role in quickly detecting prompt emission. It enables a rapid follow-up of the afterglow through the X-ray Telescope (XRT) instrument, which covers the 0.3-10 keV range (Burrows et al., 2005), and the Ultra-Violet Optical Telescope (UVOT 170 - 600 nm, (Roming et al., 2005)). Moreover, Swift’s rapid multi-wavelength afterglow follow-up has revealed new characteristics in GRB LCs (Tagliaferri et al., 2005; Nousek et al., 2006; Troja et al., 2007).
@@
Most X-ray LCs exhibit rapid decay in flux after the prompt episode, occasionally followed by flares and/or a plateau (Zhang et al., 2006; O’Brien et al., 2006; Nousek et al., 2006; Sakamoto et al., 2007; Liang et al., 2007; Willingale et al., 2007; Dainotti et al., 2008, 2010, 2016, 2017; Dereli-Bégué et al., 2024). The plateau observed in GRB LCs can be modeled using a Broken Power-Law (BPL) (Zhang et al., 2006, 2007; Racusin et al., 2009), a smooth BPL, or the phenomenological model proposed by (W07, Willingale et al. (2007)). The W07 model determines critical parameters such as the time at the end of the plateau, $T_{a}$, the corresponding flux, $F_{a}$, and the temporal index after the plateau, $\alpha_{a}$. On the other hand, the BPL model provides $T_{a}$, $F_{a}$, and the slope of the LC during the plateau, $\alpha_{1}$ and after the plateau, $\alpha_{2}$. Section  2.1 details the W07 and the parameters.
@@
The plateau phase is frequently interpreted using the magnetar model (Zhang & Mészáros, 2001; Rowlinson et al., 2014; Rea et al., 2015; Stratta et al., 2018), which attributes the emission to dipole radiations generated by the rotational energy of a newly formed Neutron Star (NS). According to this model, the plateau ends when the NS reaches its critical spin-down timescale. Uncertainties in determining $T_{a}$ are often related to the magnetar spin period and uncertainties of the magnetic field. Therefore, precise measurements of $T_{a}$ are essential to verify the validity of this model. The plateau phase, exhibiting more consistent features across various GRBs, such as length and flatness, has attracted attention because of its potential to establish relevant correlations with the plateau parameters and their application as cosmological tools. (Dainotti et al., 2008, 2010, 2011, 2013, 2015, 2017; Tang et al., 2019; Wang et al., 2022; Zhao et al., 2019; Liang et al., 2007; Li et al., 2018) have explored the luminosity at the end of the plateau, $L_{X,a}$ versus its rest-frame time $T^{*}_{X,a}$ (known as the Dainotti relation or the 2D L-T relation) $+$the rest-frame time is denoted with an asterisk. The 2D relation has also been identified in the optical plateau emissions (Dainotti et al., 2020b, 2022). Within the theoretical magnetar framework, Rowlinson et al. (2014) showed that the X-ray Dainotti relation is reproduced with a slope for $L_{a,X}$-$T^{*}_{a,X}$ of $-1$. This correlation has been applied in cosmological research, such as the development of the GRB Hubble diagram, which extends to redshifts greater than $z>8$ (Cardone et al., 2009, 2010; Postnikov et al., 2014; Dainotti et al., 2013).
@@
The 2D L-T relation has been further expanded by incorporating the peak prompt luminosity, $L_{X,peak}$, resulting in the Dainotti 3D relation (Dainotti et al., 2016, 2017, 2020a, 2022). This 3D relation has also been successfully applied to constrain cosmological parameters (Dainotti et al., 2023a, 2022g, 2022b; Cao et al., 2022a, b). Importantly, Dainotti et al. (2022g) demonstrated that reducing the uncertainties associated with the plateau emission parameters by 47.5% could achieve the same precision for the cosmological value of $\Omega_{M}$ quoted in Betoule et al. (2014) in just 8 years as we have calculated in (Dainotti et al., 2022c; Narendra et al., 2024), if we consider the optical sample and the addition of GRBs for which the redshift was inferred (Dainotti et al., 2024b, c). This improvement could be realized immediately, compared to the precision requiring 22 more years of observations under current rates and parameter uncertainties (Dainotti et al., 2022g), which highlights the significant potential of a more reliable Light Curve Reconstruction (LCR) approach, as it could substantially accelerate progress to reach the same precision achieved by SNe Ia (for details, see Dainotti et al., 2020b).
@@
In addition, GRB LCs with temporal gaps present significant challenges in testing theoretical models, such as the standard fireball model (Panaitescu & Kumar, 2000; Piran, 1999). This model is typically evaluated using closure relations (Willingale et al., 2007; Evans et al., 2009; Racusin et al., 2009; Kumar & Duran, 2010; Srinivasaragavan et al., 2020; Dainotti et al., 2021; Ryan et al., 2020; Tak et al., 2019), which involve the temporal ($\alpha$) and spectral ($\beta$) index of the afterglow. The value of $\alpha$ can correspond to $\alpha_{1}$ or $\alpha_{2}$, depending on the segment of the LC analyzed, while $\beta$ is measured for the same time window as $\alpha$. An accurate assessment of these factors is essential to evaluate the fireball model and categorize GRBs according to their morphology.
@@
Thus, here we tackle the LCR, which provides a novel approach to address the challenge of temporal gaps in LCs and it is a continuation of a previous effort by some of us (Dainotti et al., 2023b). The paper introduced a stochastic reconstruction technique using existing models and GPs, reducing uncertainties for key GRB parameters.
Recent advances in machine learning have further pushed the boundaries of LCR. For example, Demianenko et al. (2023) examined the application of neural network-based methods, including Bayesian neural networks, multilayer perceptrons, and normalization flows, to obtain approximate findings of a single LC. These methods were tested on the LCs of Supernovae type 1a taken from the Zwicky Transient Facility Bright Transient Survey and simulated PLAsTiCC. These methods demonstrated that even with limited observations, neural networks could significantly enhance the quality of the reconstruction in comparison to state-of-the-art models. Moreover, they are computationally efficient, outperforming GPs in terms of speed, and were found to be effective for subsequent tasks such as peak identification and transient classification.
@@ 2.1 Dataset and the Willingale model
This research analyzed a dataset comprising 545 GRBs. The initial sample comprises 455 GRBs detected by Swift between January 2005 and August 2019, as compiled by Srinivasaragavan et al. (2020a); Dainotti et al. (2020a). This dataset is then enlarged by additional 14 GRBs observed by Fermi-LAT, as presented in Dainotti et al. (2024a) and a further 78 GRBs detected by Swift between 2019 and 2023, as reported by Narendra et al. (2024). After accounting for two overlapping GRBs between the original and extended samples, the final dataset consists of 545 unique GRBs. The dataset includes 230 GRBs with known redshifts and 315 without redshift information, as described in Dainotti et al. (2024b, c); Narendra et al. (2024). All GRBs in the sample, including flares (about 31.5% of our sample), exhibit temporal gaps in their LCs arising from various factors. The gaps vary in size, but we fix the minimum gap size as 0.03 in the logarithmic time scale for the reconstruction task. This threshold was determined based on an analysis of all GRBs in the sample, where this size is based on the smallest possible gap. By establishing this minimum, we are realistically sampling and filling the smallest gap in the LCs.
The W07 function is to model the LCs of GRBs, defined in Eq. 1 and introduced by Willingale et al. (2007):
$f(t)=\left\{\begin{array}[]{ll}\displaystyle{F_{i}\exp{\left(\alpha_{i}\left(1-\frac{t}{T_{i}}\right)\right)}\exp{\left(-\frac{t_{i}}{t}\right)}}{\hskip 28.45274pt\rm for}\ \ t<T_{i},\\ \\ \displaystyle{F_{i}\left(\frac{t}{T_{i}}\right)^{-\alpha_{i}}\exp{\left(-\frac{t_{i}}{t}\right)}}{\hskip 28.45274pt\rm for}\ \ t\geq T_{i}.\\ \end{array}\right.$ (1)
Here, the symbols $T_{i}$ and $F_{i}$ refer to time and flux, respectively, in our case, at the end of the plateau emission, $\alpha_{i}$ is the temporal index parameter after $T_{i}$. The parameter $t_{i}$ corresponds to the onset of the rise phase, and the afterglow emission is represented by $t_{a}$. We have shown in previous papers that the onset is usually compatible with 0 (Willingale et al., 2007).
@@
The GRBs were subsequently categorized into four distinct classes based on the features of their afterglow emission, using the classification methodology established in Dainotti et al. 2023b:
@@
The first four panels of Fig.1 illustrate the LCs for the four identified classes and their corresponding W07 model fits. Our analysis relies on the analysis of GRB afterglow LCs by comparing each LC with the W07 fitted model. The parameters for each GRB LC have been derived from the work of Srinivasaragavan et al. (2020a). The blue points represent the trimmed data, where the prompt emission phase has been removed.
@@ 2.2 Reconstruction with the Willingale model
In this section, we adhere to the methodology described in Dainotti et al. (2023b) to reconstruct the LCs using the W07 model. However, our approach further extends the analysis by including all classes of GRBs, allowing for a more extensive study of the model across different categories of GRBs.
@@
We then perform the same steps described, generating histograms of the residuals for each GRB and fitting a Gaussian distribution, as a first approximation, to those residuals as presented in right bottom panel of Fig. 1. In our analysis, we followed the same methodology as outlined in Dainotti et al. (2023b), where the fitting was performed using the Gaussian, since in most of the cases the Gaussian distribution is quoted to be the best-fit in the package FindDistribution. However, there are several cases in which other distributions may have fitted better, but to provide an uniform treatment for simplicity and to better compare our analysis with the one of Dainotti et al. (2023b), we have adopted the same methodology. In future work, we will use in each case the best fit distribution for each GRB. The reconstructed flux at each time point is calculated by adding noise, sampled from the residual distribution, to the flux predicted by the W07 model:

### v5

@@ Abstract
Gamma-Ray Bursts (GRBs), observed at high-z, are probes of the evolution of the Universe and can be used as cosmological tools. Thus, we need correlations with small dispersion among key parameters. To reduce such a dispersion, we mitigate gaps in light curves (LCs), including the plateau region, key to building the two-dimensional Dainotti relation between the end time of plateau emission ($T_{a}$) and its luminosity ($L_{a}$). We reconstruct LCs using nine models: Multi-Layer Perceptron (MLP), Bi-Mamba, Fourier Transform, Gaussian Process–Random Forest Hybrid (GP-RF), Bidirectional Long Short-Term Memory (Bi-LSTM), Conditional GAN (CGAN), SARIMAX-based Kalman filter, Kolmogorov–Arnold Networks (KANs), and Attention U-Net. These methods are compared to the Willingale model (W07) over a sample of 521 GRBs. MLP and Attention U-Net outperform other methods, with MLP reducing the plateau parameter uncertainties by 37.2% for $\log T_{a}$, 38.0% for $\log F_{a}$, and 41.2% for $\alpha$ (the post-plateau slope in the W07 model), achieving the lowest 5-fold cross-validation (CV) mean squared error (MSE) of 0.0275. Attention U-Net achieved the lowest uncertainty of parameters, a 37.9% reduction in $\log T_{a}$, a 38.5% reduction in $\log F_{a}$ and a 41.4% reduction in $\alpha$, but with a higher MSE of 0.134. Although Attention U-Net achieves the largest uncertainty reduction, the MLP attains the lowest test MSE while maintaining comparable uncertainty performance, making it the more reliable model. The other methods yield MSE values ranging from 0.0339 to 0.174. These improvements in parameter precision are needed to use GRBs as standard candles, investigate theoretical models, and predict GRB redshifts through machine learning.
@@ 1 Introduction
GRBs are brief and highly luminous astrophysical phenomena detectable at remarkable distances (for review, see 2015PhR...561....1K), with observations extending up to redshift $z=9.4$ (Cucchiara et al., 2011). This property makes GRBs exceptional cosmological tools for probing the universe’s early evolution. A thorough analysis of GRBs also yields essential information about Population III stars, the first generation formed during the epoch of re-ionization. GRB emission is usually observed in two episodes: the prompt and the afterglow. The prompt phase is interpreted by internal shell collisions or magnetic reconnection and is characterized by the short duration and high energy. Hence, this is detected predominantly in $\gamma$ rays and X-rays and occasionally in optical wavelengths (Vestrand2005Natur; Blake et al., 2005; Beskin et al., 2010; 2012MNRAS.421.1874G; 2014Sci...343...38V). During the afterglow phase, which follows the primary episode, the relativistic jet impacts the circumstellar environment, transferring part of its energy. The afterglow is characterized by a long duration and energies that span a broad spectrum, including X-ray, optical, and sometimes radio bands (Costa et al., 1997; vanParadijs1997; Piro1998; Gehrels2009ARA&A).
@@
The Neil Gehrels Swift Observatory (Swift, Gehrels2004ApJ...611.1005G) is vital for detecting the temporal properties of GRBs. The Swift Burst Alert Telescope (BAT), which operates in the range of 15-150 keV (Barthelmy et al., 2005), plays a crucial role in quickly detecting prompt emission. It enables a rapid follow-up of the afterglow through the X-ray Telescope (XRT) instrument, which covers the 0.3-10 keV range (Burrows et al., 2005), and the Ultra-Violet Optical Telescope (UVOT 170 - 600 nm, (Roming2005)). Moreover, Swift’s rapid multi-wavelength afterglow follow-up has revealed new characteristics in GRB LCs (Tagliaferri2005; Nousek2006; Troja2007).
@@
Most X-ray LCs exhibit rapid decay in flux after the prompt episode, occasionally followed by flares and/or a plateau (Zhang2006; OBrien2006; Nousek2006; Sakamoto2007; Liang2007; willingale2007testing; Dainotti2008; dainotti2010a; Dainotti2016; dainotti2017a; dereli2024unraveling). The plateau observed in GRB LCs can be modeled using a Broken Power-Law (BPL) (Zhang2006; Zhang2007ApJ...655L..25Z; Racusin2009), a smooth BPL, or the phenomenological model proposed by (W07, willingale2007testing). The W07 model determines critical parameters such as the time at the end of the plateau, $T_{a}$, the corresponding flux, $F_{a}$, and the temporal index after the plateau, $\alpha_{a}$. On the other hand, the BPL model provides $T_{a}$, $F_{a}$, and the slope of the LC during the plateau, $\alpha_{1}$ and after the plateau, $\alpha_{2}$. Section  2.1 details the W07 and the parameters.
@@
The plateau phase is frequently interpreted using the magnetar model (Zhang2001; Rowlinson2014; Rea2015; Stratta2018), which attributes the emission to dipole radiations generated by the rotational energy of a newly formed Neutron Star (NS). According to this model, the plateau ends when the NS reaches its critical spin-down timescale. Uncertainties in determining $T_{a}$ are often related to the magnetar spin period and uncertainties of the magnetic field. Therefore, precise measurements of $T_{a}$ are essential to verify the validity of this model. The plateau phase, exhibiting more consistent features across various GRBs, such as length and flatness, has attracted attention because of its potential to establish relevant correlations with the plateau parameters and their application as cosmological tools. (Dainotti2008; dainotti2010a; dainotti2011a; dainotti2013determination; dainotti2015; dainotti2017a; Tang2019ApJS..245....1T; Wang:2021hcx; Zhao2019ApJ...883...97Z; Liang2007; Li2018b) have explored the luminosity at the end of the plateau, $L_{X,a}$ versus its rest-frame time $T^{*}_{X,a}$ (known as the Dainotti relation or the 2D L-T relation) +++the rest-frame time is denoted with an asterisk. The 2D relation has also been identified in the optical plateau emissions (dainotti2020b; 2022ApJS..261...25D). Within the theoretical magnetar framework, Rowlinson2014 showed that the X-ray Dainotti relation is reproduced with a slope for $L_{a,X}$-$T^{*}_{a,X}$ of 1-1. This correlation has been applied in cosmological research, such as the development of the GRB Hubble diagram, which extends to redshifts greater than $z>8$ (Cardone et al., 2009, 2010; postnikov2014nonparametric; dainotti2013).
@@
The 2D L-T relation has been further expanded by incorporating the peak prompt luminosity, $L_{X,peak}$, resulting in the Dainotti 3D relation (Dainotti2016; dainotti2017a; dainotti2020a; 2022ApJS..261...25D). This 3D relation has also been successfully applied to constrain cosmological parameters (dainotti2023b; dainotti2022g; dainotti2022b; Cao et al., 2022a, b). Importantly, dainotti2022g demonstrated that reducing the uncertainties associated with the plateau emission parameters by 47.5% could achieve the same precision for the cosmological value of $\Omega_{M}$ quoted in Betoule et al. (2014) in just 8 years as we have calculated in (dainotti2022c; narendra2024grb), if we consider the optical sample and the addition of GRBs for which the redshift was inferred (giovanna2024inferring; dainotti2024optical). This improvement could be realized immediately, compared to the precision requiring 22 more years of observations under current rates and parameter uncertainties (dainotti2022g), which highlights the significant potential of a more reliable Light Curve Reconstruction (LCR) approach, as it could substantially accelerate progress to reach the same precision achieved by SNe Ia (for details, see dainotti2020x).
@@
In addition, GRB LCs with temporal gaps present significant challenges in testing theoretical models, such as the standard fireball model (panaitescu2000analytic; piran1999gamma). This model is typically evaluated using closure relations (willingale2007testing; evans2009methods; racusin2009jet; kumar2010external; srinivasaragavan2020investigation; dainotti2021closure; ryan2020gamma; tak2019closure), which involve the temporal ($\alpha$) and spectral ($\beta$) index of the afterglow. The value of $\alpha$ can correspond to $\alpha_{1}$ or $\alpha_{2}$, depending on the segment of the LC analyzed, while $\beta$ is measured for the same time window as $\alpha$. An accurate assessment of these factors is essential to evaluate the fireball model and categorize GRBs according to their morphology.
@@
Thus, here we tackle the LCR, which provides a novel approach to address the challenge of temporal gaps in LCs and it is a continuation of a previous effort by some of us (dainotti2023stochastic). The paper introduced a stochastic reconstruction technique using existing models and GPs, reducing uncertainties for key GRB parameters.
Recent advances in machine learning have further pushed the boundaries of LCR. For example, demianenko2023understanding examined the application of neural network-based methods, including Bayesian neural networks, multilayer perceptrons, and normalization flows, to obtain approximate findings of a single LC. These methods were tested on the LCs of Supernovae type 1a taken from the Zwicky Transient Facility Bright Transient Survey and simulated PLAsTiCC. These methods demonstrated that even with limited observations, neural networks could significantly enhance the quality of the reconstruction in comparison to state-of-the-art models. Moreover, they are computationally efficient, outperforming GPs in terms of speed, and were found to be effective for subsequent tasks such as peak identification and transient classification.
@@ 2.1 Dataset and the Willingale model
This research analyzed a dataset comprising 521 GRBs. The initial sample comprises 455 GRBs detected by Swift between January 2005 and August 2019, as compiled by 2020ApJ...903...18S; dainotti2020a. This dataset is then enlarged by additional 14 GRBs observed by Fermi-LAT, as presented in dainotti2024analysis and a further 78 GRBs detected by Swift between 2019 and 2023, as reported by narendra2024grb. After accounting for two overlapping GRBs between the original and extended samples, the final dataset consists of 521 unique GRBs. The dataset includes 230 GRBs with known redshifts and 315 without redshift information, as described in giovanna2024inferring; dainotti2024optical; narendra2024grb. All GRBs in the sample, including flares (about 31.5% of our sample), exhibit temporal gaps in their LCs arising from various factors. To address irregular gaps in GRB LCs, we implemented a gap-aware reconstruction in log-time space. A minimum gap threshold of 0.05 was imposed. The number of reconstructed points was adaptively set as a fraction of the original LC size: 5% for LCs with more than 500 points, 10% for those with 250–500 points, 30% for 100–250 points, and 40% for fewer than 100 points (minimum of 20 points). Extremely sparse LCs (about 5% of the sample with <<26 points) with very large gaps, reconstruction was not feasible. Thus, the dataset was reduced to 521 GRBs.
The W07 function is to model the LCs of GRBs, defined in Eq. 1 and introduced by willingale2007testing:
[…]
Here, the symbols $T_{i}$ and $F_{i}$ refer to time and flux, respectively, in our case, at the end of the plateau emission, $\alpha_{i}$ is the temporal index parameter after $T_{i}$. The parameter $t_{i}$ corresponds to the onset of the rise phase, and the afterglow emission is represented by $t_{a}$. We have shown in previous papers that the onset is usually compatible with 0 (willingale2007testing).
@@
The GRBs were subsequently categorized into four distinct classes based on the features of their afterglow emission, using the classification methodology established in dainotti2023stochastic:
@@
The first four panels of Fig.1 illustrate the LCs for the four identified classes and their corresponding W07 model fits. Our analysis relies on the analysis of GRB afterglow LCs by comparing each LC with the W07 fitted model. The parameters for each GRB LC have been derived from the work of 2020ApJ...903...18S. The blue points represent the trimmed data, where the prompt emission phase has been removed.
@@ 2.2 Reconstruction with the Willingale model
In this section, we adhere to the methodology described in dainotti2023stochastic to reconstruct the LCs using the W07 model. However, our approach further extends the analysis by including all classes of GRBs, allowing for a more extensive study of the model across different categories of GRBs.
@@
We then perform the same steps described, generating histograms of the residuals for each GRB and fitting a Gaussian distribution, as a first approximation, to those residuals as presented in right bottom panel of Fig. 1. In our analysis, we followed the same methodology as outlined in dainotti2023stochastic, where the fitting was performed using the Gaussian, since in most of the cases the Gaussian distribution is quoted to be the best-fit in the package FindDistribution. However, there are several cases in which other distributions may have fitted better, but to provide an uniform treatment for simplicity and to better compare our analysis with the one of dainotti2023stochastic, we have adopted the same methodology. In future work, we will use in each case the best fit distribution for each GRB. The reconstructed flux at each time point is calculated by adding noise, sampled from the residual distribution, to the flux predicted by the W07 model: